OpenStudy (anonymous):

Solve 4x + 4(6y) + 8z = 140(6)

OpenStudy (angela210793):

4x+24y+8z=140*6???? solve it how?

OpenStudy (anonymous):

It was a word problem and that was the equation i got out of it.

OpenStudy (anonymous):

Pastaroni is a local Italian restaurant that serves a specialized pasta salad. Included in this salad is macaroni, linguine and bow-tie pasta. The salad contains six times as much linguine as it does macaroni. Both macaroni and linguine sell for $4; bow-tie pasta sells for $8. If the restaurant makes 140 pounds of pasta every day, how many pounds of each type of pasta are needed if the salad sells for $6?

OpenStudy (radar):

Pssssst you are pretty good at expressing your problems in a manner that confuses, frustrates, and otherwise discourages any help by others. Above is a good example.

OpenStudy (radar):

Let x = number of pounds of macroni then 6x = number of pounds of linguine then 140-6x-x = number of pounds of bow-tie pasta or 140-7x. Now do the cost. $4x = cost of macaroni $24x = cost lingjine and $8(140-7x) = cost of bow tie. total cost is 140 times $6 or $840.00 thus the formula will be: $4x+$24x +$8(140+7)=$840 now solve for x $28x+$1120-$56x=$840 -28x=-280 x=10 or ten pounds of macaroni 6x=60 or 60 pounds of linguine Then that means 70 pounds of bow tie pasta. Following assumption made (and a very unrealistic one) he was selling the salad at cost!!

OpenStudy (anonymous):

I Need Help With This One , So Is This The Answer @radar & @Pssssst